# Projective limits of probabilistic symmetries and their applications to random graph limits

### v1

@@ 2.1 Point processes
Let $\mathbb{X}$ be a Polish space, with topology $T$ and denote by $\mathcal{X}:=\sigma(T)$ the associated Borel $\sigma$-algebra, and let $M(\mathbb{X})$ to be the set of locally finite simple counting measures on $\mathbb{X}$.
We now endow $M(\mathbb{X})$ with a measurable structure. For this we define for every $X\in\mathcal{X}$ and $\xi\in M(\mathbb{X})$ the evaluation map $\epsilon_{X}:\xi\mapsto\xi(X)$. We then let $\mathcal{T}_{\epsilon}$ denote the smallest topology such that the maps $\epsilon_{X}$ are continuous. The associated $\sigma$-algebra is denoted by $\mathcal{M}:=\sigma(\mathcal{T}_{\epsilon})$. This is the smallest $\sigma$-algebra that makes the evaluation maps $\epsilon_{X}$ measurable. We note that the described construction is the standard way to endow the space of locally finite simple counting measures on $\mathbb{X}$ with a topology and measurable structure.
@@ 2.2 Projective limits
Here we will briefly discuss the notions of projective limits of topological spaces and of measures on them.
@@ 2.2.1 Projective limits of topological spaces
The first property states that each smaller space $\mathbb{X}_{n}$ can be obtained as a “projection” of the limit space $\mathbb{X}_{\mathbb{N}}$, in a manner that respects the local projections $\pi_{mn}$. The second property is referred to as the universal property. We usually write $\langle\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}},\pi_{n}\rangle=\underleftarrow{\mathrm{lim}}\,\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ to denote that $\langle\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}},\pi_{n}\rangle$ is the projective limit of the system $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$.
@@ 2.2.2 Projective limits of probability measures
The notion of projective limits of topological spaces $(\mathbb{X}_{n},T_{n})$ extends naturally to projective limits of probability measures on the corresponding measurable spaces with sigma-algebras $\sigma(T_{n})$.
@@
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of topological spaces. A collection $(\mu_{n})_{n\geq 1}$ of probability measures on $\mathbb{X}_{n}$ is called a projective system of measures if $\mu_{n}=\pi_{mn}\ast\mu_{m}$ for all $n<m$, i.e., if the following diagram commutes for all $n<m$:
@@
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of topological spaces and $(\mu_{n})_{n\geq 1}$ be a projective sequence of probability measures with respect to $\pi_{mn}$. Then the projective limit of $(\mu_{n})_{n\geq 1}$ (if it exists) is the unique probability measure $\mu_{\mathbb{N}}$ on $\mathbb{X}_{\mathbb{N}}$ such that $\mu_{n}=\pi_{n}\ast\mu_{\mathbb{N}}$ holds for all $n\geq 1$, where $\pi_{n}$ are the projections $\mathbb{X}_{\mathbb{N}}\to\mathbb{X}_{n}$:
@@
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of topological spaces and let $(\mu_{n})_{n\geq 1}$ be a projective sequence of probability measures on $(\mathbb{X}_{n},\sigma(T_{n}))$ with respect to $\pi_{mn}$. Then there exist a unique probability measure $\mu_{\mathbb{N}}$ on $(\mathbb{X}_{\mathbb{N}},\sigma(T_{\mathbb{N}}))$ such that $\mu_{n}=\pi_{n}\ast\mu_{\mathbb{N}}$ for all $n\geq 1$, where $\pi_{n}$ are the canonical projections of the projective limit space.
@@ 2.3 Direct limits
Our results will couple projective limits of probability measures to direct limits of groups acting on them. Direct limits are in a certain sense inverse to projective limits, which are also known as inverse limits. We refer to [15, Chapter 7] or [39, Chapter 3] for a general treatment, while here we use direct limits in application to groups.
@@
    $\Phi_{n}$$\Phi_{m}$$\Phi_{\mathbb{N}}$$\scriptstyle{\eta_{nn}}$$\scriptstyle{\eta_{n}}$$\scriptstyle{\eta_{m}}$ (3)
@@
Let $\langle\Phi_{n},\eta_{nm}\rangle_{\mathbb{N}}$ be a direct system of groups, denote by $\sqcup_{n\geq 1}\Phi_{n}$ the disjoint union of $\Phi_{n}$, and define on it the equivalence relation $\sim$ by
$\varphi_{n}\sim\varphi_{m}\text{ if and only if }\eta_{kn}(\varphi_{n})=\eta_{km}(\varphi_{m})$
with $k=\max\{n,m\}$. Now define the group
@@
with group operation $\varphi_{n}][\varphi_{m}]=[\eta_{kn}(\varphi_{m})\eta_{km}(\varphi_{m})]$ with $k=\max\{n,m\}$. Define also the homomorphisms $\eta_{n}:\Phi_{n}\to\Phi_{\mathbb{N}}$ by $\eta_{n}(\varphi_{n})=[\varphi_{n}]$. Then
@@
$\Phi_{n}$$\Phi_{m}$$\Phi_{\infty}$$\scriptstyle{\eta_{nn}}$$\scriptstyle{\eta_{n}}$$\scriptstyle{\eta_{m}}$ (5)
2.4 Limits of probabilistic symmetries
@@
We next add another ingredient to the picture, a projective system of probability measures. Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ and $\langle\Phi_{n},\eta_{nm}\rangle_{\mathbb{N}}$ be a compatible system of topological spaces and groups, and let, in addition, $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be equipped with a projective system of probability measures as in Definition 2.3. Then it follows immediately from Definition 2.9 that
@@
3 Main results: existence and symmetry of projective limits of point processes
Our main results are about projective limits of point processes and their invariance. Specifically, we identify the symmetry group with respect to which the projective limit of a point process is invariant.
Yet we start with a general result, showing that projective limits of probability measures that are invariant with respect to a compatible direct system of groups, is invariant with respect to the direct limit of these groups.
@@ 3 Main results: existence and symmetry of projective limits of point processes
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ and $\langle\Phi_{n},\eta_{nm}\rangle_{\mathbb{N}}$ be a compatible projective system of topological spaces and direct system of groups acting on $\mathbb{X}_{n}$. Furthermore, let $\langle\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}},\pi_{n}\rangle$ and $\langle\Phi_{\mathbb{N}},\eta_{n}\rangle$ denote, respectively, their projective and direct limits. Then the group $\Phi_{\mathbb{N}}$ acts on $(\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}})$ compatibly with respect to $\pi_{n}$, i.e., the diagram (8) commutes. Moreover, if $(\mu_{n})_{n\geq 1}$ is a projective system of probability measures on $(\mathbb{X}_{n},\sigma(T_{n}))$ that are invariant with respect to $\Phi_{n}$, then the projective limit $\mu_{\mathbb{N}}=\underleftarrow{\mathrm{lim}}\,\mu_{n}$ is invariant with respect to $\Phi_{\mathbb{N}}=\underrightarrow{\mathrm{lim}}\,\Phi_{n}$.
Consider now a projective system $(\mu_{n})_{n\geq 1}$ of point processes on $\mathbb{X}_{n}$ that are invariant under a direct system of groups. Then it follows from Theorem 3.1 that the projective limit of these point processes $\mu_{\mathbb{N}}$, which is a probability measure on the projective limit of the $\langle M(\mathbb{X}_{n}),\sigma(T_{n}),\pi_{mn}\rangle_{\mathbb{N}}$, is invariant under the direct limit of the groups. This is nice, but the main problem with this observation is that according to Theorem 3.1, the projective limit $\mu_{\mathbb{N}}$ is a general probability measure on a projective limit space, while we actually want $\mu_{\mathbb{N}}$ to be a point process on a given topological space. For this to be the case, we need to impose some additional conditions that would allow us to show that the projective limit of $\langle M(\mathbb{X}_{n}),\sigma(T_{n}),\pi_{mn}\rangle_{\mathbb{N}}$ is a space of locally finite measures $M(\mathbb{X}_{\infty})$ on some space $\mathbb{X}_{\infty}$.
@@
To define the projective structure we denote by $\iota_{nm}:\mathbb{X}_{n}\hookrightarrow\mathbb{X}_{m}$ and $\iota_{n}:\mathbb{X}_{n}\hookrightarrow\mathbb{X}$ the canonical inclusions, and note that these are continuous. Then we can define the projections $\pi_{mn}:M(\mathbb{X}_{m})\to M(\mathbb{X}_{n})$ by
$\pi_{mn}\xi=\xi\circ\iota_{nm},$ (9)
where $\circ$’ denotes the standard composition of functions. Similarly, we define the projections $\pi_{n}:M(\mathbb{X})\to M(\mathbb{X}_{n})$ by
$\pi_{n}\xi=\xi\circ\iota_{n}.$ (10)
@@
However, this result is not sufficiently strong for some applications, in particular those that we will consider in the next section. For such applications, we want to go beyond the direct limit group because this group turns out to be not the full limit group with respect to which we want to show invariance of our limit point process. The direct limit group is instead a dense finitary subgroup of this full group. To illustrate what we mean, let $\Phi_{n}$ be the group of permutation on $n]:=\{1,2,\dots,n\}$. They form a direct system whose direct limit $\Phi_{\mathbb{N}}$ is the group of permutations of $\mathbb{N}$ that only change a finite number of elements [31, Chapter 6]. However, we really want to show the invariance of our limit measure with respect to the group $\Phi_{\infty}$ of all permutations on $\mathbb{N}$, and this is where our definition of a compatible direct pre-limit group comes into play.
Yet for technical reasons, we need one more restriction on the connection between a projective limit probability measure $\mu_{\mathbb{N}}$ on $M(\mathbb{X})$ and its pre-limit symmetry group $\Phi_{\infty}$.
Definition 3.1 (Finite characteristic).
Let $(\Phi_{n})_{n\geq 1}$ be a directed system of groups, and $\Phi_{\infty}$ be a direct pre-limit. The invariance of a probability measure $\mu$ on $M(\mathbb{X})$ with respect to $\Phi_{\infty}$ has a finite characteristic if there exists a family of subsets $\Psi_{n}\subseteq\Phi_{n}$, $n\geq 1$, such that the invariance of $\mu$ with respect to $\Phi_{\infty}$ holds if and only if
$\mu(A)=\mu((\eta_{n}\varphi_{n})^{-1}A),\quad\forall n\geq 1,\,\varphi_{n}\in\Psi_{n},\text{ and }A\in\mathcal{X}.$
With this last definition in place, we are ready to formulate our last main result.
@@
Let $(\mu_{n})_{n\geq 1}$ be a projective system of point processes on $\mathbb{X}_{n}$ with respect to $\pi_{mn}$, as in Theorem 3.2, and let $\mu_{\mathbb{N}}=\underleftarrow{\mathrm{lim}}\,\mu_{n}$ be its projective limit. Let $(\Phi_{n})_{n\geq 1}$ be a compatible directed system of groups acting on $\mathbb{X}_{n}$, and let $\Phi_{\infty}$ be a direct pre-limit group acting on $\mathbb{X}$, whose invariance has a finite characteristic with respect to $\mu_{\mathbb{N}}$. If $\mu_{n}$ is invariant with respect to $\Phi_{n}$ for every $n\geq 1$, then $\mu_{\mathbb{N}}$ is invariant with respect to $\Phi_{\infty}$.
@@
4 Applications to random graphs

### v2

@@ 2.1 Point processes
Let $\mathbb{X}$ be a Polish space, with topology $T$ and denote by $\mathcal{X}:=\sigma(T)$ the associated Borel $\sigma$-algebra, and let $M(\mathbb{X})$ be the set of locally finite simple counting measures on $\mathbb{X}$.
We now endow $M(\mathbb{X})$ with its canonical measurable structure. For every $X\in\mathcal{X}$ and $\xi\in M(\mathbb{X})$ we define the evaluation map $\epsilon_{X}:M(\mathbb{X})\to\mathbb{N}\cup\{\infty\}$ by $\epsilon_{X}(\xi):=\xi(X)$, and let $\mathcal{M}$ be the $\sigma$-algebra generated by these maps, $\mathcal{M}:=\sigma\bigl(\epsilon_{X}:X\in\mathcal{X}\bigr)$. This is the standard $\sigma$-algebra on the space of locally finite simple counting measures used in the theory of point processes [37]. We recall that $\mathcal{M}$ is equal to the Borel $\sigma$-algebra generated by the vague topology $\mathcal{T}$ on $M(\mathbb{X})$, i.e. the topology generated by the integration maps $\pi_{f}:\xi\mapsto\int f\mathrm{d}\xi$ for $f\in C_{c}(\mathbb{X})$ which is the set of continuous functions $f$ on $\mathbb{X}$ with bounded support [31, Theorem 4.2].
@@ 2.2 Projective limits
Here we will briefly discuss the notions of projective limits of topological spaces and of measures on the corresponding Borel spaces. We consider topological spaces, and not general measurable spaces, as our space of interest $(M(\mathbb{X}),\mathcal{M})$ is a Borel space with respect to the vague topology and this is an essential ingredient for studying limits of probability measures.
@@ 2.2.1 Projective limits of topological spaces
The first property states that each smaller space $\mathbb{X}_{n}$ can be obtained as a “projection” of the limit space $\mathbb{X}_{\mathbb{N}}$, in a manner that respects the local projections $\pi_{mn}$. The second property is referred to as the universal property.
We usually write $\langle\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}},\pi_{n}\rangle=\underleftarrow{\mathrm{lim}}\,\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ to denote that $\langle\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}},\pi_{n}\rangle$ is the projective limit of the system $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$.
@@ 2.2.2 Projective limits of probability measures
The notion of projective limits of topological spaces $(\mathbb{X}_{n},T_{n})$ extends naturally to projective limits of probability measures. Here we assume that these spaces are Polish spaces equipped with their Borel sigma-algebras $\sigma(T_{n})$ and probability measures on them.
@@
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of Polish spaces. A collection $(\mu_{n})_{n\geq 1}$ of probability measures on $\mathbb{X}_{n}$ is called a projective system of measures if $\mu_{n}=\pi_{mn}\ast\mu_{m}$ for all $n<m$, i.e., if the following diagram commutes for all $n<m$:
@@
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of Polish spaces and $(\mu_{n})_{n\geq 1}$ be a projective sequence of probability measures with respect to $\pi_{mn}$. Then the projective limit of $(\mu_{n})_{n\geq 1}$ (if it exists) is the unique probability measure $\mu_{\mathbb{N}}$ on $\mathbb{X}_{\mathbb{N}}$ such that $\mu_{n}=\pi_{n}\ast\mu_{\mathbb{N}}$ holds for all $n\geq 1$, where $\pi_{n}$ are the projections $\mathbb{X}_{\mathbb{N}}\to\mathbb{X}_{n}$:
@@
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be a projective system of Polish spaces and let $(\mu_{n})_{n\geq 1}$ be a projective sequence of probability measures on $(\mathbb{X}_{n},\sigma(T_{n}))$ with respect to $\pi_{mn}$. Then there exist a unique probability measure $\mu_{\mathbb{N}}$ on $(\mathbb{X}_{\mathbb{N}},\sigma(T_{\mathbb{N}}))$ such that $\mu_{n}=\pi_{n}\ast\mu_{\mathbb{N}}$ for all $n\geq 1$, where $\pi_{n}$ are the canonical projections of the projective limit space.
@@ 2.3 Direct limits
Our results will couple projective limits of probability measures to direct limits of groups acting on them. Direct limits are in a certain sense inverse to projective limits, which are also known as inverse limits. We refer to [15, Chapter 7] or [40, Chapter 3] for a general treatment, while here we use direct limits in application to groups.
@@
    $\Phi_{n}$$\Phi_{m}$$\Phi_{\mathbb{N}}$$\scriptstyle{\eta_{nm}}$$\scriptstyle{\eta_{n}}$$\scriptstyle{\eta_{m}}$ (3)
@@
Let $\langle\Phi_{n},\eta_{nm}\rangle_{\mathbb{N}}$ be a direct system of groups, denote by $\sqcup_{n\geq 1}\Phi_{n}$ the disjoint union of $\Phi_{n}$, and define on it the equivalence relation $\sim$ between $\varphi_{n}\in\Phi_{n}$ and $\varphi_{m}\in\Phi_{m}$ by
[…]
Now define the group
@@
with group operation $\varphi_{n}][\varphi_{m}]=[\eta_{nk}(\varphi_{n})\eta_{mk}(\varphi_{m})]$ with some $k\geq\max\{n,m\}$. Define also the homomorphisms $\eta_{n}:\Phi_{n}\to\Phi_{\mathbb{N}}$ by $\eta_{n}(\varphi_{n})=[\varphi_{n}]$. Then
@@
$\Phi_{n}$$\Phi_{m}$$\Phi_{\infty}$$\scriptstyle{\eta_{nm}}$$\scriptstyle{\eta_{n}}$$\scriptstyle{\eta_{m}}$ (5)
2.4 Compatible limits
@@
We next add another ingredient to the picture, a projective system of probability measures. Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ and $\langle\Phi_{n},\eta_{nm}\rangle_{\mathbb{N}}$ be a compatible system of Polish spaces and groups, and let, in addition, $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ be equipped with a projective system of probability measures as in Definition 2.3. Then it follows immediately from Definition 2.9 that
@@
3 Main results: limits of probabilistic symmetries
We start with a general result, showing that projective limits of probability measures that are invariant with respect to a compatible direct system of groups, is invariant with respect to the direct limit of these groups.
@@ 3 Main results: limits of probabilistic symmetries
Let $\langle\mathbb{X}_{n},T_{n},\pi_{mn}\rangle_{\mathbb{N}}$ and $\langle\Phi_{n},\eta_{nm}\rangle_{\mathbb{N}}$ be a compatible projective system of Polish spaces and direct system of groups acting on $\mathbb{X}_{n}$. Furthermore, let $\langle\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}},\pi_{n}\rangle$ and $\langle\Phi_{\mathbb{N}},\eta_{n}\rangle$ denote, respectively, their projective and direct limits. Then the group $\Phi_{\mathbb{N}}$ acts on $(\mathbb{X}_{\mathbb{N}},T_{\mathbb{N}})$ compatibly with respect to $\pi_{n}$, i.e., the diagram (8) commutes. Moreover, if $(\mu_{n})_{n\geq 1}$ is a projective system of probability measures on $(\mathbb{X}_{n},\sigma(T_{n}))$ that are invariant with respect to $\Phi_{n}$, then the projective limit $\mu_{\mathbb{N}}=\underleftarrow{\mathrm{lim}}\,\mu_{n}$ is invariant with respect to $\Phi_{\mathbb{N}}=\underrightarrow{\mathrm{lim}}\,\Phi_{n}$.
The rest of our main results are about projective limits of point processes and their invariance. Specifically, we identify the symmetry group with respect to which the projective limit of a point process is invariant.
Consider a projective system $(\mu_{n})_{n\geq 1}$ of point processes on $\mathbb{X}_{n}$ that are invariant under a direct system of groups. Then it follows from Theorem 3.1 that the projective limit of these point processes $\mu_{\mathbb{N}}$, which is a probability measure on the projective limit of the $\langle M(\mathbb{X}_{n}),\sigma(T_{n}),\pi_{mn}\rangle_{\mathbb{N}}$, is invariant under the direct limit of the groups. This is nice, but the main problem with this observation is that according to Theorem 3.1, the projective limit $\mu_{\mathbb{N}}$ is a general probability measure on a projective limit space, while we actually want $\mu_{\mathbb{N}}$ to be a point process on a given Polish space. For this to be the case, we need to impose some additional conditions that would allow us to show that the projective limit of $\langle M(\mathbb{X}_{n}),\sigma(T_{n}),\pi_{mn}\rangle_{\mathbb{N}}$ is a space of locally finite measures $M(\mathbb{X}_{\infty})$ on some space $\mathbb{X}_{\infty}$.
@@
To define the projective structure we denote by $\iota_{nm}:\mathbb{X}_{n}\hookrightarrow\mathbb{X}_{m}$ and $\iota_{n}:\mathbb{X}_{n}\hookrightarrow\mathbb{X}$ the canonical inclusions, and note that these are continuous. Then we can define the projections $\pi_{mn}:M(\mathbb{X}_{m})\to M(\mathbb{X}_{n})$, as a pull back, by
$\pi_{mn}\xi(A_{n})=\xi(\iota_{nm}(A_{n})),$ (9)
for $A_{n}$ a measurable set in $M(\mathbb{X}_{n})$. Similarly, we define the projections $\pi_{n}:M(\mathbb{X})\to M(\mathbb{X}_{n})$ by
$\pi_{n}\xi(A)=\xi(\iota_{n}(A_{n})).$ (10)
@@
However, this result is not sufficiently strong for some applications, in particular those that we will consider in the next section. For such applications, we want to go beyond the direct limit group because this group turns out to be not the full limit group with respect to which we want to show invariance of our limit point process. The direct limit group is instead a dense finitary subgroup of this full group. To illustrate what we mean, let $\Phi_{n}$ be the group of permutation on $n]:=\{1,2,\dots,n\}$. They form a direct system whose direct limit $\Phi_{\mathbb{N}}$ is the group of permutations of $\mathbb{N}$ that only change a finite number of elements [32, Chapter 6]. However, we really want to show the invariance of our limit measure with respect to the group $\Phi_{\infty}$ of all permutations on $\mathbb{N}$, and this is where our definition of a compatible direct pre-limit group comes into play.
@@
Let $(\mu_{n})_{n\geq 1}$ be a projective system of point processes on $\mathbb{X}_{n}$ with respect to $\pi_{mn}$, as in Theorem 3.2, and let $\mu_{\mathbb{N}}=\underleftarrow{\mathrm{lim}}\,\mu_{n}$ be its projective limit. Let $(\Phi_{n})_{n\geq 1}$ be a compatible directed system of groups acting on $\mathbb{X}_{n}$, and let $\Phi_{\infty}$ be a direct pre-limit group acting on $\mathbb{X}$, such that any $\mu$ on $M(\mathbb{X})$ is invariant with respect to $\Phi_{\infty}$ if
$\mu(A)=\mu((\eta_{n}\varphi_{n})^{-1}A),\quad\forall n\geq 1,\,\varphi_{n}\in\Phi_{n},\text{ and }A\in\mathcal{X}.$
If $\mu_{n}$ is invariant with respect to $\Phi_{n}$ for every $n\geq 1$, then $\mu_{\mathbb{N}}$ is invariant with respect to $\Phi_{\infty}$.
@@
4 Applications to limits of random graphs